The area of a rhombus is 143m2. If the longer diagonal is 26m,
find the length of one side of the rhombus and its altitude.
1
Expert's answer
2020-09-27T16:37:53-0400

The area of any convex quadrilateral can be determined as "S = \frac{1}{2}d_1 d_2 \sin\varphi" , where "d_1" and "d_2" are the diagonals, "\varphi" is the angle of diagonals' intersection (90 degrees for a rhombus).

Thus, the formula simplifies to "S = \frac{1}{2}d_1 d_2". From here, "d_2 = \frac{2S}{d_1}=11" m.

Then, if we take one quarter of a rhombus, we will have a right triangle with diagonals' halves as shorter sides of triangle and side of the rhombus as hypotenuse.

Hence, side of the rhombus a = "\sqrt{13^2+5.5^2}=14.11559\approx14.12" m. <---------------------------

The last thing to calculate is the altitude h. From "S=a\cdot h" we get "h=\frac{S}{a}=10.13064\approx10.13" m. <-----------------------------